# Leveraging Priors via Diffusion Bridge for Time Series Generation

### v1

@@ Introduction
Diffusion Prior Design with TimeBridge
To address these questions, we first explore suitable prior designs for diffusion synthesis. Afterward, we propose TimeBridge, a general diffusion bridge model that leverages the aforementioned prior selections, elaborating its application in both unconditional and conditional tasks.
Data- and Time-dependent Priors for Time Series
@@
Data-Dependent Prior.
@@
To investigate RQ1, we suggest using data prior to having a data-dependent distribution to better approximate the data distribution. Sampling from $\mathcal{N}(\mathbf{0},\mathbf{I})$ lacks data-specific information, which can be enhanced by enforcing that the priors capture data scale and temporal dependencies. Recently, the usages of data-dependent prior have been investigated for audio (Popov et al. 2021; Lee et al. 2022) and image (Yu et al. 2024) domains. Therefore, we test the use of data-dependent prior for time series by setting the prior distribution as follows:
${\mathbf{x}}_{T}\sim\mathcal{N}(\bm{\mu},\texttt{diag}(\bm{\sigma}^{2})),$ (10)
@@
where $\bm{\mu}$ and $\bm{\sigma}^{2}$ are independently calculated in the same dimension of the time series ${\mathbf{x}}$ in $\mathbb{R}^{\tau\times d}$. $\texttt{diag}(\bm{\sigma}^{2})$ indicates the diagonal covariance matrix of each element.
Time-Dependent Prior.
@@
As a response to RQ2, we focus on capturing temporal dependencies in noise addition towards prior distribution. Recently, Biloš et al. (2023) injected stochastic processes instead of random noise to preserve the continuity of temporal data and Han, Zheng, and Zhou (2022) utilized the pre-trained model for prediction and set the mean of prior as the output for regression.
As a solution, we employ Gaussian processes (GPs) for prior selection, which is effective for modeling temporal data (Biloš et al. 2023; Ansari et al. 2024). GPs are distributions over functions with time input $k$ characterized by the mean function $\bm{m}(k)$ and the positive definite kernel $\mathcal{K}(k,k^{\prime})$. We use the radial basis function (RBF) kernel, defined as $\mathcal{K}(k,k^{\prime})=\eta\exp(-\gamma|k-k^{\prime}|^{2})$, to ensure a stationary process. To construct a time-dependent prior based on data-dependent properties, we add a variance term to the kernel function to encode data information into the prior distribution. Consequently, we construct the prior incorporating of the temporal dynamics as follows:
${\bm{x}}_{T}\sim\mathcal{GP}(\bm{m},\mathcal{K}+\bm{\Sigma}),$ (11)
where $\bm{\Sigma}=[\bm{\Sigma}_{ij}]$ is a correlation matrix function in which $\bm{\Sigma}_{ij}$ represents the correlation of the data between timestamps $i$ and $j$. For simplified version of Equation (10), we set $\bm{m}$ as $\bm{\mu}$ and $\bm{\Sigma}$ as $\texttt{diag}(\bm{\sigma}^{2})$ to align with the data-dependent prior, i.e., ${\mathbf{x}}_{T}\sim\mathcal{N}(\bm{\mu},\mathcal{K}+\texttt{diag}(\bm{\sigma}^{2}))$.
Scale-preserving Sampling for Conditional Priors
@@
We now consider the condition ${\mathbf{y}}\in\mathcal{X}$, provided as a time series, which is the most prevalent case for real-world settings. For example, a given trend can guide the synthesis output or fixed points at specific time steps can determine the properties of synthetic data. We explore the use of a prior on the same data scale instead of relying solely on conditional embedding in standard diffusion.
Soft constraints with given trends.
@@
To answer RQ3, we argue that the diffusion bridge is well-suited for preserving a given condition by setting the prior as the condition. The diffusion bridge has shown its effectiveness in conditional tasks such as coloring (Zhou et al. 2024), image-to-image translation (Li et al. 2023; Liu et al. 2023), and improving naive base predictions (Chen et al. 2023b; Lee et al. 2024).
@@
Conditional generation based on trend is straightforward by (i) setting the pair-wise condition to prior ${\mathbf{x}}_{T}={\mathbf{y}}$ and (ii) training the translation from trend to data. As the expected mean of ${\mathbf{x}}_{t}$ is a linear combination of ${\mathbf{x}}_{0}$ and ${\mathbf{x}}_{T}={\mathbf{y}}$ in Equation (6), the model learns the correction starting from the trend samples, utilizing the same data scale. In contrast, standard diffusion models are unsuitable for translating conditions into synthetic data; instead, they generate new data based on a specific trend condition, typically requiring additional steps such as providing guidance or correcting intermediate data samples ${\mathbf{x}}_{t}$ during sampling. Thus, we can also eliminate the need for additional penalty functions or value correction during sampling.
Hard constraints with fixed points.
@@
In RQ4, preserving data points of hard constraints can be important in conditional tasks. To ensure the fixed point condition during sampling, we introduce a novel point-preserving sampling method that prevents adding noise to the values to be fixed. This is a unique trait of our framework because standard diffusion models cannot preserve identity trajectories to the constrained values from Gaussian noise. We make details of the proposed sampling method in Appendix C.
@@
The direct case of fixed point constraints is imputation. With the mask ${\mathbf{m}}$ indicating the missing value as 0, we construct the condition ${\mathbf{c}}={\mathbf{y}}\odot{\mathbf{m}}$ for imputation. However, unlike soft constraints that apply to all values in ${\mathbf{y}}\in\mathbb{R}^{\tau\times d}$, the condition for imputation ${\mathbf{c}}\in\mathbb{R}^{\tau\times d}$ has missing values, which should be interpolated to data-dependent values rather than zeros. Therefore, we build local linear models for condition-to-time series imputation. We adopt linear spline interpolation (De Boor 1978) to generate a prior for fixed point constraints. In a feature-wise manner, let us consider $x\in\mathbb{R}$ to be a single dimensional input of ${\bm{x}}\in\mathbb{R}^{d}$. With a set $K$ of feature-wise observed times for $x$, the prior $x_{T}$ with the interpolation of each feature with timestamp $k$ is:
$x_{T}(k)=x(k_{j-1})+\frac{x(k_{j})-x(k_{j-1})}{k_{j}-k_{j-1}}(k-k_{j-1}),$ (12)
@@
Note that before the first observation, $k<k_{0}$, $x_{T}(k)=x(k_{0})$. After the last observation, $k\geq k_{|K|-1}$, $x_{T}(k)=x(k_{|K|-1})$. The detailed Algorithm is in Appendix C.
TimeBridge: Bridge Modeling for Time Series
Based on the aforementioned prior selections, we propose TimeBridge, a general diffusion bridge model for time series synthesis. Our approach yields high-quality samples with fewer sampling steps and is compatible with existing time series diffusion frameworks.
For score matching in Equation (9), we build the denoiser $D_{\theta}$ for time series. For architecture design, we adopt the backbone of Diffusion-TS (Yuan and Qiao 2024), i.e., Fourier-based loss function (Zhou et al. 2022; Liu et al. 2024; Nguyen et al. 2022; Yi et al. 2024) and transformer encoder-decoder models with seasonal-trend decomposition (Zeng et al. 2023; Wu et al. 2021). With the bridge diffusion framework, our objective function is formulated as follows:
$\mathcal{L}_{\theta}=\mathbb{E}_{t,{\mathbf{x}}_{0}}[w_{t}(\lVert{\mathbf{x}}_{0}-D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})\rVert^{2}$ (13)
$w_{t}$ indicates the weight scheduler of the loss function and $\lambda$ indicates the strength of fast Fourier transform (FFT). We construct the bridge model to produce output by decomposing the input into trend and seasonal components as follows:
$D_{\theta}=V_{t}^{tr}(\theta,{\mathbf{x}}_{t},t,{\mathbf{x}}_{T})+\sum_{i=1}^{K}S_{i,t}(\theta,{\mathbf{x}}_{t},t,{\mathbf{x}}_{T})+R(\theta,{\mathbf{x}}_{t},t,{\mathbf{x}}_{T}),$ (14)
where $V_{t}^{tr}$ is the output for the trend synthesis layer, $S_{i,t}$ is the output for each seasonal synthesis layer $i$ among $K$ layers and $R$ is the output for estimated residual.
As time series data possess stochasticity (Shen and Kwok 2023), we directly predict $D_{\theta}$ rather than using parametrization technique from noise prediction used in other domains (Karras et al. 2022; Zhou et al. 2024). Moreover, we utilize the second-order Heun sampler (Ascher and Petzold 1998; Karras et al. 2022), achieving both quality and efficiency. Overall, our TimeBridge model leverages the flexibility of prior selection as a diffusion bridge designed for time series.
@@
The baseline experimental results (denoted by $\dagger$) are adopted from (Yuan and Qiao 2024). We re-implement Diffusion-TS with their official code.
Table 1: Quality comparison of unconditional time series synthetic data. The best results are in bold and the second best are underlined. We report the average score (lower is better) and rank of each method.
Related Works
Standard diffusion models for time series.
Time series diffusion models have become the de facto method in synthesis tasks. Rasul et al. (2021) initially used diffusion networks for time series based on recurrent networks. For imputation, SSSD (Alcaraz and Strodthoff 2023) and CSDI (Tashiro et al. 2021) considered time series imputation similar to image inpainting tasks. TimeDiff (Shen and Kwok 2023), LDT (Feng et al. 2024), and TMDM (Li et al. 2024) focused on time series forecasting tasks. Recently, Coletta et al. (2024) investigated guiding the sampling with a penalty function called DiffTime. Diffusion-TS (Yuan and Qiao 2024) introduced a decomposition architecture to disentangle time series data into trend and seasonal components and applied a Fourier-based loss term. They also expanded their unconditional model to conditional generation tasks by guiding the diffusion sampling toward the input conditions.
Existing time series bridge models.
Recently, some researchers have investigated the use of the diffusion bridge in the time series domain. Chen et al. (2023a) investigated the convergence analysis of the Schrödinger bridge algorithm and applied them to time series imputation tasks. Hamdouche, Henry-Labordere, and Pham (2023) used kernel estimation for the Schrödinger bridge combined with feed-forward and LSTM networks, especially for the application of deep hedging on real-data sets. Garg, Zhang, and Zhou (2024) relaxed the Schrödinger bridge by setting a geometric mixture of targets and other distributions for the prior distribution and applied their method to time series. However, most of them used iterative solutions for bridge problems, which required a heavy computational burden and incompatibility with recent standard diffusion models. Furthermore, none of them suggested a general framework for both unconditional and conditional time series generation.
Experiments
Experimental Setup
@@
For measures, we assess four metrics for unconditional time series generation on a normalized (0-1) scale. Context-Fréchet Inception Distance (Context-FID) score (Jeha et al. 2022) measures the distributional quality for mean and variance projected on embedding on trained TS2Vec (Yue et al. 2022). Correlational score (Liao et al. 2020) measures temporal dependency of real and synthetic data using the absolute error of cross-correlation. Discriminative score (Yoon, Jarrett, and Van der Schaar 2019) measures the performance of a classifier trained to distinguish original and synthetic data in a supervised manner. Additionally, for imputation tasks, we calculate the mean squared error (MSE) and mean absolute error (MAE) on each data scale.
Our experiments are conducted using PyTorch on four NVIDIA GeForce RTX 4090 GPUs. We report the mean and its interval of five experiments. For the detailed settings, refer to Appendix C.
(a) Sines
(b) Mujoco
(c) ETTh
(d) Stocks
(e) Energy
(f) fMRI
Figure 3: Visualization on t-SNE. Red and blue indicate original and synthetic data, respectively.
@@
Correlation score Trend Baseline 0.0493±.0104 0.0372±.0139 0.0524±.0194 0.4816±.0633 0.4816±.0999 0.5246±.0915 0.2711 (-)
@@
Table 2: Quality comparison of trend-conditioned time series synthetic data.
Figure 4: Illustration of sampling path from (Left) prior to (Right) data samples on the ETTh.
@@ Methods
The baseline experimental results are adopted from (Alcaraz and Strodthoff 2023) and (Yuan and Qiao 2024).
Table 3: Imputation results of MSE in the order of 1e-3 on the Mujoco dataset with a random mask of ratio {0.7, 0.8, 0.9}. The best results are in bold and the second best are underlined.
@@
Table 4: Imputation result of MSE and MAE on the ETTh and Energy datasets with a geometric mask of ratio {0.25, 0.5, 0.75}. The best results are in bold and the second best are underlined.
Unconditional Generation
@@
We demonstrate the unconditional generation tasks with a length of 24 for the aforementioned datasets and evaluate the performance of the previous measures. For baseline methods, we compare with TimeVAE (Desai et al. 2021), TimeGAN (Yoon, Jarrett, and Van der Schaar 2019), Cot-GAN (Xu et al. 2020), Diffwave (Kong et al. 2021), DiffTime (Coletta et al. 2024) with SSSD (Alcaraz and Strodthoff 2023) backbone, and Diffusion-TS (Yuan and Qiao 2024). For the proposed TimeBridge, we choose the data-dependent settings in Equation (10) and TimeBridge-GP denotes the time-dependent GP prior in Equation (11).
@@
Results are shown in Table 1. We observe that the proposed method outperforms the previous methods. The performance gain is significant in context-FID and discriminative scores, reducing the scores of 40.70% and 18.05% on average, respectively. Notably, the strength of data and time-dependent prior is shown in real datasets, such as ETTh and Stocks. This indicates that our prior selection and modeling help in approximating the data distribution during synthesis.
@@
To visualize the generated data, we plot the t-SNE plots with the generated data in Figure 3. The red original data and the blue synthetic data show the overlapped plot in almost all datasets. The generated samples well approximate the distribution properties in complex datasets, such as the Stocks and Energy datasets. For further visualization concerning PCA and kernel density plots, refer to Appendix E.
Synthesis with Trend Condition

### v2

@@ 4. Prior Design for Time Series Synthesis
4.1. Data- and Time-dependent Priors for Time Series
@@
Data-Dependent Prior
@@
To investigate RQ1, we suggest using data prior to having a data-dependent distribution to better approximate the data distribution. Sampling from $\mathcal{N}(\mathbf{0},\mathbf{I})$ lacks data-specific information, which can be enhanced by enforcing that the priors capture data scale and temporal dependencies. Recently, the usages of data-dependent prior have been investigated for audio (Popov et al., 2021; Lee et al., 2022) and image (Yu et al., 2024) domains. Therefore, we test the use of data-dependent prior for time series by setting the prior distribution as follows:
(19) ${\mathbf{x}}_{T}\sim\mathcal{N}(\bm{\mu},\texttt{diag}(\bm{\sigma}^{2})),$
@@
where $\bm{\mu}$ and $\bm{\sigma}^{2}$ are the mean and variance of the data samples for each feature and timestamp independently calculated in the same dimension of the time series (e.g., for ETTh, resulting in $\bm{\mu}$ with a shape of $(24\times 7)$). $\texttt{diag}(\bm{\sigma}^{2})$ indicates the diagonal covariance matrix of each element.
Time-Dependent Prior
@@
In response to RQ2, we focus on capturing temporal dependencies in noise addition towards prior distribution. Recently, Biloš et al. (2023) injected stochastic processes instead of random noise to preserve the continuity of temporal data and Han et al. (2022) utilized the pre-trained model for prediction and set the mean of prior as the output for regression.
As a solution, we employ Gaussian processes (GPs) for prior selection, which is effective for modeling temporal data (Biloš et al., 2023; Ansari et al., 2024). GP is a distribution over function that defines the joint variability of an arbitrary pair of time points. For given sequence of timestamps $\Omega=\{1,\ldots,\tau\}$ and a function $\bm{f}_{GP}$ of timestamps that follows a GP with mean function $\bm{m}(\cdot)$ and the covariance function $\bm{\Sigma}(\cdot)$, we can draw ${\mathbf{x}}_{T}$ as follows:
(20) $\bm{f}_{GP}\sim{GP}(\bm{m},\bm{\Sigma}),\ \ {\mathbf{x}}_{T}\sim\bm{f}_{GP}(\Omega).$
In this paper, we construct the time-dependent prior by aligning the GP with the aforementioned data-dependent prior at given time points. Specifically, we add the radial basis function (RBF) kernel $\mathcal{K}(i,j)=\eta exp(-\gamma|i-j|^{2})$ for timestamps $i$ and $j$ into Equation (19). Details for sampling method are in Appendix B.4.
4.2. Scale-preserving for Conditional Priors
@@
We now consider the condition ${\mathbf{y}}\in\mathcal{X}$, provided as a time series, which is the most prevalent case for real-world settings. We explore using a prior on the same data scale instead of relying solely on conditional embeddings in standard diffusion.
Soft constraints with given trends
@@
To answer RQ3, we argue that the diffusion bridge is well-suited for preserving a given condition by setting the prior as the condition. The diffusion bridge has shown its effectiveness in conditional tasks such as coloring (Zhou et al., 2024), image-to-image translation (Li et al., 2023; Liu et al., 2023), and improving naive base predictions (Chen et al., 2023b; Lee et al., 2024).
@@
Conditional generation based on trend is straightforward by (i) setting the pair-wise condition to prior ${\mathbf{x}}_{T}={\mathbf{y}}$ and (ii) training the translation from trend to data. As the expected mean of ${\mathbf{x}}_{t}$ is a linear combination of ${\mathbf{x}}_{0}$ and ${\mathbf{x}}_{T}={\mathbf{y}}$ in Equation (9), the model learns the correction starting from the trend samples, utilizing the same data scale. In contrast, standard diffusion models are unsuitable for translating conditions into synthetic data; instead, they generate new data based on a specific trend condition, typically requiring additional steps such as providing guidance or correcting intermediate data samples ${\mathbf{x}}_{t}$ during sampling. Thus, we can also eliminate the need for penalty functions or corrections for sampling.
Hard constraints with fixed points
@@
In RQ4, preserving data points of hard constraints can be important in conditional tasks. To ensure the fixed point condition during sampling, we introduce a novel point-preserving sampling method that prevents adding noise to the values to be fixed. This is a unique trait of our framework because standard diffusion models cannot preserve identity trajectories to the constrained values from Gaussian noise.
@@
The direct case of fixed point constraints is imputation. With the mask ${\mathbf{m}}$ indicating the missing value as 0, we construct the condition ${\mathbf{c}}={\mathbf{y}}\odot{\mathbf{m}}$ for imputation where $\odot$ indicates the Hadamard product. However, unlike soft constraints that apply to all values in ${\mathbf{y}}\in\mathbb{R}^{\tau\times d}$, the condition for imputation ${\mathbf{c}}\in\mathbb{R}^{\tau\times d}$ has missing values, which should be interpolated to data-dependent values rather than zeros. Therefore, we build local linear models for condition-to-time series imputation. We adopt linear spline interpolation (De Boor, 1978) to generate a prior for fixed point constraints. In a feature-wise manner, let us consider $x\in\mathbb{R}$ as a single-dimensional input of ${\bm{x}}\in\mathbb{R}^{d}$. With a set $K$ of feature-wise observed times for $x$, the prior $x_{T}$ with the interpolation of each feature with timestamp $k$ is:
(21) $x_{T}(k)=x(k_{j-1})+\frac{x(k_{j})-x(k_{j-1})}{k_{j}-k_{j-1}}(k-k_{j-1}),$
@@
Note that before the first observation, $k<k_{0}$, $x_{T}(k)=x(k_{0})$. After the last observation, $k\geq k_{|K|-1}$, $x_{T}(k)=x(k_{|K|-1})$. We make details of the point-preserving sampler and its usage in imputation in Algorithm 1.
5. Related Works
Time Series Generation Models
Diffusion models show better performance in synthesis tasks, compared to time series generation models with VAE (Naiman et al., 2024), GANs (Yoon et al., 2019). Rasul et al. (2021) initially used diffusion networks for time series based on recurrent networks. For imputation, SSSD (Alcaraz and Strodthoff, 2023) and CSDI (Tashiro et al., 2021) considered time series imputation similar to image inpainting tasks. TimeDiff (Shen and Kwok, 2023), LDT (Feng et al., 2024), and TMDM (Li et al., 2024) focused on time series forecasting tasks. Coletta et al. (2024) investigated guiding the sampling with a penalty function called DiffTime. Our network design follows Diffusion-TS (Yuan and Qiao, 2024), which introduced a decomposition architecture to disentangle time series data into trend and seasonal components and applied a Fourier-based loss term. Recent studies also explore frequency-domain diffusion (Crabbé et al., 2024) and transformer-based alternatives (Chen et al., 2024).
Schrödinger Bridge in Time Series
Several studies have brought Schrödinger bridges to time series tasks. Chen et al. (2023a) analyzed algorithmic convergence and applied the bridge to imputation. Hamdouche et al. (2023) combined kernel-based bridges with feed-forward and LSTM networks for deep hedging. Garg et al. (2024) relaxed the Schrödinger bridge by setting a geometric mixture for the prior distribution and applied their method to time series. All of these methods require iterative solvers, which are computationally expensive and incompatible with standard diffusion pipelines. Also, None of them provides a unified framework for both unconditional and conditional generation. To the best of our knowledge, our paper is the first to use diffusion bridge, extending beyond classical Schrödinger bridge formulations and providing flexible priors.
Input: Model $D_{\theta}({\mathbf{x}}_{t})$, score function ${\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i}},t_{i},{\mathbf{x}}_{T},{T})$, diffusion steps $\{0=t_{0}<\cdots<t_{\Gamma}=T\}$, Condition ${\mathbf{y}}$, masking ${\mathbf{m}}$, step ratio $s$
Output: Denoised output ${\mathbf{x}}_{0}$
if Imputation  then ${\mathbf{x}}_{T}\leftarrow$ Linear spline interpolation using Equation (21) with masking ${\mathbf{m}}$;
for $i=\Gamma,\ldots,1$ do
       Sample $\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0,I})$
       $\tilde{t}_{i}\leftarrow t_{i}+s(t_{i-1}-t_{i})$
       if $i\neq 1$ then
Algorithm 1 Point-preserving Sampler for TimeBridge
Table 2. Quality comparison of unconditional time series synthetic data with length of 24. The best results are in bold and the second best are underlined. We report the average score (lower is better) and the rank of each method. The baseline experimental results (denoted by $\dagger$) are adopted from (Yuan and Qiao, 2024). We re-implement Diffusion-TS with their official code.
@@ 5. Related Works
Table 3. Quality comparison of unconditional time-series synthetic data with a longer length of 64. Lower is better.
Dataset Method C.-FID ($\downarrow$) Corr. ($\downarrow$) Disc. ($\downarrow$)
ETTh Diffusion-TS 0.235±.011 0.061±.010 0.078±.017
TimeBridge-GP 0.110±.009 0.038±.011 0.045±.013
Stocks Diffusion-TS 0.948±.227 0.005±.005 0.149±.035
TimeBridge-GP 0.693±.165 0.005±.003 0.090±.018
Energy Diffusion-TS 0.067±.010 0.442±.077 0.065±.011
TimeBridge-GP 0.072±.010 0.548±.102 0.149±.005
fMRI Diffusion-TS 3.662±.104 4.585±.068 0.245±.234
TimeBridge-GP 0.527±.024 1.569±.028 0.211±.184
Average Diffusion-TS 1.228 1.273 0.134
TimeBridge-GP 0.351 0.540 0.124
6. Experiments
6.1. Experimental Setup
@@
For measures, we assess three metrics for unconditional time series generation on a normalized (0-1) scale. Context-Fréchet Inception Distance (Context-FID, C.-FID) score (Jeha et al., 2022) measures the distributional quality for mean and variance projected on embedding on trained TS2Vec (Yue et al., 2022). Correlational (Corr.) score (Liao et al., 2020) measures temporal dependency of real and synthetic data using the absolute error of cross-correlation. Discriminative (Disc.) score (Yoon et al., 2019) measures the performance of a classifier trained to distinguish original and synthetic data in a supervised manner. For imputation, we use the mean squared error (MSE) and mean absolute error (MAE).
Our experiments are conducted using PyTorch on four NVIDIA GeForce RTX 4090 GPUs. We report the mean and its interval of five experiments. For the detailed settings, refer to Appendix B.
Table 4. Quality comparison of trend-conditioned time series synthetic data.
@@
Correlational score Trend Baseline 0.0493±.0104 0.0372±.0139 0.0524±.0194 0.4816±.0633 0.4816±.0999 0.5246±.0915 0.2711 (-)
@@
Table 5. Imputation results of MSE in the order of 1e-3 on the Mujoco dataset with a random mask of ratio {0.7, 0.8, 0.9}. The best results are in bold and the second best are underlined. The baseline experimental results are adopted from (Alcaraz and Strodthoff, 2023) and (Yuan and Qiao, 2024).
@@ Methods
Table 6. Imputation result of MSE and MAE on the ETTh and Energy datasets with a geometric mask of ratio {0.25, 0.5, 0.75}. The best results are in bold and the second best are underlined.
@@
6.2. Unconditional Generation
@@
We demonstrate the unconditional generation tasks with a length of 24 as baselines. We compare with TimeVAE (Desai et al., 2021), TimeGAN (Yoon et al., 2019), Cot-GAN (Xu et al., 2020), Diffwave (Kong et al., 2021), DiffTime (Coletta et al., 2024) with SSSD (Alcaraz and Strodthoff, 2023) backbone, and Diffusion-TS (Yuan and Qiao, 2024). For the proposed TimeBridge, we choose the data-dependent settings in Equation (19) and TimeBridge-GP denotes the time-dependent GP prior in Equation (20). As the prior mean and variance are derived from the training datasets, the proposed method does not require additional information.
@@
Results are shown in Table 2. We observe that the proposed method outperforms the previous methods. The performance gain is significant in Context-FID and discriminative scores, reducing the scores of 40.70% and 18.05% on average, respectively. Notably, the strength of data and time-dependent prior is shown in real datasets, such as ETTh and Stocks. This indicates that our prior selection and modeling help in approximating the data distribution during synthesis. In average rank, the proposed method shows the best performance. It indicates the power of prior selection in time series generation within a similar structure.
We compare our model with Diffusion-TS on long-horizon generation tasks with a sequence length of 64. The results in Table 3 show that, on average, TimeBridge-GP produces higher-quality samples than standard Gaussian. Additional experiments with even longer sequences are reported in Appendix C.
@@
For visualization, we show the t-SNE plots with the generated data in Figure 3. The red original data and the blue synthetic data show an overlapped plot in almost all datasets. The generated samples well-approximate the distribution properties in complex datasets, such as the Stocks and Energy datasets.
(a) Sines
(b) Mujoco
(c) ETTh
(d) Stocks
(e) Energy
(f) fMRI
Figure 3. Visualization on t-SNE. Red and blue indicate original and synthetic data, respectively.
6.3. Synthesis with Trend Condition